### v1

@@ Abstract
Accurate and efficient seismic response prediction is essential for the design of resilient structures. While the Finite Element Method (FEM) remains the standard for nonlinear seismic analysis, its high computational demands limit its scalability and real-time applicability. Recent developments in deep learning—particularly Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and Long Short-Term Memory (LSTM) models—have shown promise in reducing the computational cost of the nonlinear seismic analysis of structures. However, these data-driven models often struggle to generalize and capture the underlying physics, leading to reduced reliability. We propose a novel Physics-Informed U-Net-LSTM framework that integrates physical laws with deep learning to enhance both accuracy and efficiency. By embedding domain-specific constraints into the learning process, the proposed model achieves improved predictive performance over conventional Machine Learning (ML) architectures. This hybrid approach bridges the gap between purely data-driven methods and physics-based modeling, offering a robust and computationally efficient alternative for seismic response prediction of structures.
keywords:
[…]
{graphicalabstract}[Uncaptioned image]{highlights}
A novel PhyULSTM framework is developed for data-driven structural seismic modeling.
Physics laws in loss guide learning, prevent overfitting, and improve generalization.
The model can predict seismic response without physical constraints as well.
@@ 1 Introduction
The increasing number and intensity of natural disasters such as earthquakes, typhoons, and tsunamis have heightened the global demand for resilient and adaptive infrastructure systems. Among these hazards, seismic events pose particularly severe threats to structural integrity, leading to widespread damage and loss of life. Accurate prediction of structural response under seismic loading is thus a cornerstone of resilient design. Conventional numerical methods, including the Newmark-$\beta$ [newmark1959method], KR-$\alpha$ [kolay2014development], and the finite element method (FEM)[reddy1993introduction, huebner2001finite, zienkiewicz2005finite], have long been the foundation for modeling dynamic behavior of structures. These techniques are reliable and grounded in physics, yet they are computationally intensive, especially when applied to nonlinear analysis of a full-scale structure. Their high computational cost limits real-time applications and scalability.
@@
In recent years, machine learning (ML) and artificial intelligence (AI) have emerged as promising alternatives or supplements to traditional physics-based modeling. A range of ML techniques,including support vector machines (SVM) [yinfeng2008nonlinear, segura2020metamodel, gharehbaghi2020estimating], multilayer perceptrons (MLP) [NN_ref1, NN_ref2, NN_ref3, NN_ref4, NN_ref5, NN_ref6, NN_ref7, NN_ref8, NN_ref9, lagaros2012neural], deep neural networks [vaidyanathan2005artificial, kim2019response, kim2020probabilistic], long short-term memory networks (LSTMs) [zhang2019deep], and convolutional neural networks (CNNs) [wu2019deep], have been employed to develop surrogate models for rapid prediction of structural response. These approaches reduce computational cost while offering acceptable levels of accuracy. RNNs and LSTMs, in particular, are often adopted for modeling temporal dynamics in structural systems due to their ability to learn long-term dependencies in time series data [graves2012long]. Meanwhile, CNNs with dilated filters have shown promise for time-series modeling by enabling access to long-range dependencies without recurrence [bai2018empirical, oord2016wavenet]. These models have also been successfully applied in domains such as ECG classification, structural health monitoring, continuum damage mechanics and anomaly detection [bird2020cross, tang2019convolutional, danilczyk2021smart, zhao2019speech, pantidis2023integrated, pantidis2024fenn].
@@
Building upon these advances, physics-informed machine learning has gained traction within earthquake engineering and structural dynamics community, where it addresses critical challenges such as data scarcity and limited model generalization. By embedding governing physical laws directly into ML architectures, physics-informed frameworks enhance predictive accuracy and ensure physical consistency. For instance, physics-informed LSTM frameworks (e.g., PhyLSTM2, PhyLSTM3)[phylstm] estimate key variables like displacement and restoring force by incorporating these laws into the loss function during training. Similarly, PhyCNN [phycnn] integrates domain knowledge with deep CNNs to predict nonlinear seismic responses through a data-driven approach. These innovative methods have been applied across diverse scenarios, including subway station dynamics [huang2021deep], vehicle–bridge interactions [li2021dynamic], and dam behavior prediction using CNN–GRU architectures [li2022new]. More recently, PI-LSTM has been introduced and benchmarked against PhyCNN and PhyLSTM, further demonstrating the potential of physics-guided ML architectures for structural response prediction [liu2023pi]. Recent works [raissi2018deep, raissi2019physics, sun2020surrogate, zhu2019physics] have highlighted the importance of incorporating physical constraints—such as ordinary and partial differential equations and boundary conditions—directly into deep learning models to improve stability and accuracy even when data is limited.
Despite the progress made, key limitations persist. LSTMs and RNNs, while effective in modeling temporal dependencies, suffer from long training times and are prone to overfitting, particularly in low-data regimes [bengio1994learning, trinh2018learning]. Deep CNNs, although powerful, require substantial computational resources when applied to long time series, and their performance degrades under large plastic deformations [kuo2024gnn]. Moreover, capturing long-range interactions and nonlinearities in dynamic systems remains a fundamental challenge. To overcome these challenges, we propose a novel Physics-Informed U-net–LSTM architecture for the data-driven modeling of structural seismic response. The model incorporates a causal one-dimensional U-Net, developed leveraging the “slow feature” hypothesis [wiskott2002slow], to extract hierarchical temporal features across multiple scales while preserving temporal causality. These features are then passed to an LSTM network, enabling efficient sequence modeling and dynamic prediction. Importantly, physical knowledge—expressed through governing equations of motion—is embedded into the loss function, guiding the model to honor the laws of mechanics even in data-scarce scenarios. The framework is validated on both numerical and experimental benchmarks and is compared with PhyCNN framework. Even without knowledge of the structural system, it can still predict the response only by utilizing the data loss in a purely data-driven manner. Results demonstrate that the proposed approach offers superior accuracy, robustness, and generalization, especially under conditions of strong nonlinearity, making it a viable surrogate for seismic response prediction of complex structural systems.
@@ 2 Physics-Informed U-net-LSTM Network
In this section, we present the PhyULSTM framework for surrogate modeling of nonlinear dynamic structural systems subjected to seismic ground motion. PhyULSTM consists of three core components: a 1D U-Net, a deep LSTM network, and a graph-based tensor differentiator. The proposed 1D U-Net, a tailored variant of the standard U-Net architecture, is designed to work synergistically with deep LSTM networks for time-series analysis. It has potential to alleviate common limitations of traditional RNNs and LSTMs—such as difficulty in capturing long-range temporal dependencies and susceptibility to overfitting—especially when processing high-resolution seismic inputs like ground acceleration. This is achieved by leveraging the ‘Slow Feature’ hypothesis [wiskott2002slow], which posits that the most informative features in temporal sequences evolve slowly over time, even when raw signals exhibit rapid fluctuations. By extracting such temporally stable features across multiple time scales and applying causal convolutions—which restrict each prediction to depend only on current and past inputs—the U-Net enhances the LSTM’s ability to learn long-term dependencies within an auto-regressive framework. Additionally, physical knowledge is explicitly incorporated into the model through a graph-based tensor differentiator (based on the finite difference method), as introduced by Zhang et al. [phycnn] in their PhyCNN framework. This component computes derivatives of the state-space outputs to construct a physics-informed loss term derived from the system’s governing equations. The hybrid U-Net–LSTM model is trained on seismic input–output datasets—obtained either through numerical simulation or sensor measurements—and subsequently deployed as a surrogate model for predicting structural responses under future seismic excitations. A detailed illustration of the PhyULSTM framework is presented in Figure 1.
@@
where $M$ is the mass matrix; $x(t)$ and $\dot{x}(t)$ denote the relative displacement and velocity with respect to the ground; $\ddot{x}_{g}(t)$ is the ground acceleration vector; $\Gamma$ represents the force distribution vector; and $h(t)$ denotes the latent generalized restoring force vector, which is modeled as a function of the ground acceleration $\ddot{x}_{g}(t)$ and the structural acceleration with respect to the ground $\ddot{x}(t)$. Normalizing Eq. (1) by $M$, the governing equation can be rewritten as
@@
where $g(t)=M^{-1}h(t)$ is the mass-normalized restoring force. The explicit form of $g(t)$ is assumed unknown and is therefore learned implicitly by the proposed PhyULSTM framework.
Figure 1: The proposed Physics-Informed U-net-LSTM Network (PhyULSTM) for time-series modeling. The 1D U-net first receives the ground acceleration data as input and extracts features at multiple time scales and adapts it to our auto-regressive scenario by ensuring that the convolutions are causal. The outputs from the U-net are then fed to the deep LSTM network that maps the temporal feature maps to the corresponding output space. The outputs are state space variables z(t){z}(t) = $\{x(t),\dot{x}(t),g(t)\}$. Available physics knowledge is incorporated directly into the loss function. In addition to it, a graph-based tensor differentiator using the central finite difference method developed by Zhang et.al in their PhyCNN framework is implemented to calculate the derivative of state space outputs ${z_{t}}(t)$ = $\{x_{t}(t),\dot{x}_{t}(t),g_{t}(t)\}$ to construct the physics loss from the governing equation, where the subscript t represents the derivative of the state with respect to time. By optimizing the network hyperparameters $\theta=\{W_{\theta},b_{\theta}\}$, PhyULSTM can interpret the measurement data (e.g. $\{x_{m},\dot{x}_{m},g_{m}\}$) While satisfying the physical equation of motion in equation 2, e.g. $f\rightarrow 0$. Here $W_{\theta}$$b_{\theta}$ are weights and biases of the neural network. $J_{D}(\theta)$ is the data loss based on the measurements and $J_{P}(\theta)$ denotes the physics loss, which imposes a physical constraint on the neural network.
@@
Figure 2: Proposed 1D version of U-net: The U-Net architecture is mainly divided into four major components: the encoder blocks, bottleneck, decoder blocks, and the output convolution block. The encoder blocks perform convolution and pooling operations, generating skip connections and providing input to subsequent encoder blocks after pooling. The decoder blocks involve upsampling, followed by concatenation with the skip connections from the corresponding encoder block outputs, and subsequent convolution operations. The bottleneck layer, serving as a bridge, consists solely of a convolution block, linking the encoder and decoder paths. The output from the final decoder block is fed to the output convolution block, where it first passes through a Conv1D layer with a number of filters equal to the number of channels required in the final output (kernel size = 1), and sigmoid activation, followed by a second Conv1D layer and a linear activation function, which further processes the output to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension. Here, ‘ConvBlock’is a fundamental building block consisting of a convolutional layer, followed by Batch Normalisation, and a ReLU activation function.
U-Net [ronneberger2015u] is an established framework originally developed for biomedical image segmentation [unet1, unet2]. It has a symmetrical U-shaped structure, consisting of an encoder and a decoder connected by skip connections. The encoder captures contextual information by progressively downsampling the input image, while the decoder reconstructs the image by upsampling and combining high-resolution features from the encoder, ensuring precise localisation.
@@
The ‘Slow Feature Analysis’ hypothesis [wiskott2002slow] posits that meaningful features in dynamic signals often evolve slowly over time. Considering the encoder-decoder components and the ability to capture contextual information and combine resolution features, U-nets are already doing it, i.e. computing features at different time scales with two-dimensional convolutions and combining them to make predictions at the same resolution as the input. One such adaptation is the one-dimensional U-Net, suitable for time-series analysis [stoller2019sequnetonedimensionalcausalunet]. In this study, we propose a causal 1D U-Net to process seismic time-series data that captures such features across multiple temporal scales while maintaining causal structure, ensuring outputs at a given time step depend only on current and past inputs. This enhances the model’s suitability for sequential tasks and integration with LSTM networks.
@@
The architecture comprises four parts: the encoder (contractive path), the bottleneck, the decoder part (expansive path) and finally, the output convolution block. The input is processed sequentially through two encoder blocks, having 50 and 100 filters respectively. Each encoder block generates two outputs: one is a skip connection (or shortcut connection) for the respective decoder block, and the other (output after 1D pooling operation) serves as the input for the subsequent encoder block. The final encoder block provides inputs for the bottleneck layer and also creates a skip connection. The encoder block comprises two sequential ‘ConvBlock’ units. Each ‘ConvBlock’ executes a series of operations: a 1D convolution ( kernel size = 2, padding = ‘causal’ ), followed by Batch Normalization, and finally a ReLU activation function.
The resulting output from this process serves as both a skip connection (shortcut path) and as the input for the subsequent MaxPooling1D operation. MaxPooling1D is a down-sampling operation typically used in convolutional neural networks (CNNs). This method reduces the dimensionality of the input tensor by extracting the maximum value from within a sliding window, known as the pool size (set to 2), across the input. So, the encoder block consists of two main components: the convolution blocks and the max pooling block.
@@
For down-sampling, the ‘Conv1D’ operation with strides=2 (padding= ‘causal’) can also be implemented, but here in this case ‘MaxPooling1D’ is used to emphasize keeping feature extraction and down-sampling as distinct operations. And to ensure that the most significant features (e.g. sudden large significant peaks in ground acceleration data) are preserved during downsampling. There are also a few additional advantages. It is more straightforward than implementing ‘Conv1D’ with a stride of 2, and it reduces feature map size without adding extra learnable parameters.
The bottleneck layer also comprises two sequential ’ConvBlock’ units, using 200 filters without any pooling operation. This layer acts as a crucial bridge between the encoder and decoder, capturing complex, high-level features while maintaining the spatial dimensions of the feature maps. In the decoder stage, the first decoder block receives the output from the bottleneck layer, along with the skip connection from the second encoder block, and uses these inputs to upsample the feature map.
@@
Within the decoder block, a 1D upsampling operation is performed on the input tensor that repeats each temporal step size times along the time axis. In this case, as the parameter size=2 is specified, the layer performs upsampling by repeating each temporal step of the input sequence twice along the time axis. This step helps to gradually reconstruct the original input size. After the 1D upsampling operation, the feature map is concatenated with the corresponding skip connection from the encoder blocks. Then the output passes through two ConvBlocks. This sequence is repeated in each of the two decoder blocks, with 100 and 50 filters, respectively, to progressively refine the feature maps and reconstruct the final output with high accuracy and detail. The output from the final decoder block is then first passed through a Conv1D layer with a number of filters equal to the number of channels required in the final output, a kernel size of 1, and sigmoid activation, which reduces the channels to match the dimension of the final output required and maps the output values between 0 and 1. The output has the same length as the input due to padding. This is followed by a second Conv1D layer and a linear activation function, which further processes the output by expanding the channels to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension, this is very crucial otherwise the deep LSTM network cannot be implemented subsequently. The network architecture is shown in Figure 2 and represented in the table 1. The U-Net output feeds directly into the LSTM module.
Layer Type Number of Filters Kernel Size Stride Activation
@@
Output Convolutional Number of output channels 1 1 Sigmoid
Final output Convolutional Number of output channels 1 1 Linear
Table 1: Proposed 1D U-Net Architecture. Note: Inside each Conv Block (Convolution Block), a consecutive set of the following operations: Convolution → Batch Normalization → ReLU activation takes place.

### v2

@@ Abstract
Accurate and efficient seismic response prediction is essential for the design of resilient structures. While the Finite Element Method (FEM) remains the standard for nonlinear seismic analysis, its high computational demands limit its scalability and real-time applicability. Recent developments in deep learning—particularly Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), and Long Short-Term Memory (LSTM) models have shown promise in reducing the computational cost of the nonlinear seismic analysis of structures. However, these data-driven models often struggle to generalize and capture the underlying physics, leading to reduced reliability. We propose a novel Physics-Informed U-Net-LSTM framework that integrates physical laws with deep learning to enhance both accuracy and efficiency. The proposed 1D U-net captures the underlying latent features of the long-term input sequences. And by embedding domain-specific constraints into the learning process, the proposed model achieves improved predictive performance over conventional Machine Learning (ML) architectures. This approach bridges the gap between purely data-driven methods and physics-based modeling, offering a robust and computationally efficient alternative for predicting the seismic response of structures.
Keywords Nonlinear structural dynamics, Seismic response, Earthquake engineering, PhyULSTM, LSTM, U-net, CNNs
@@ 1 Introduction
The increasing number and intensity of natural disasters such as earthquakes, typhoons, and tsunamis have heightened the global demand for resilient and adaptive infrastructure systems. Among these hazards, seismic events pose particularly severe threats to structural integrity, often resulting in widespread damage and significant loss of life. To ensure structural safety and reliability, and to limit potential damage, accurate prediction of structural response under seismic loading is essential. Conventional numerical methods, such as the Newmark-$\beta$ method [32], KR-$\alpha$ [20], and the finite element method (FEM)[39, 16, 57], have long been employed to model the dynamic behavior of structures. These techniques are reliable and firmly grounded in physical principles; however, they are computationally intensive, particularly when applied to nonlinear analysis of full-scale structures. The high computational cost associated with these methods limits their suitability for real-time applications and large-scale parametric studies.
@@
In recent years, machine learning (ML) and artificial intelligence (AI) have emerged as promising alternatives or supplements to traditional numerical methods. A range of ML techniques, including support vector machines (SVM) [50, 41, 10], multilayer perceptrons (MLP) [13, 5, 4, 51, 17, 36, 14, 30, 24, 23], deep neural networks [47, 18, 19], long short-term memory networks (LSTMs) [52], and convolutional neural networks (CNNs) [49, 25, 26, 27, 21], have been employed to develop surrogate models for the prediction of the seismic response of structures. These approaches reduce computational cost while offering acceptable levels of accuracy. RNNs and LSTMs, in particular, are often adopted for modeling temporal dynamics in structural systems due to their ability to learn long-term dependencies in time series data [11]. Whereas, CNNs with dilated filters have shown promise for time-series modeling by enabling access to long-range dependencies without recurrence [1, 33]. These models have also been successfully applied in domains such as ECG classification, structural health monitoring, continuum damage mechanics and anomaly detection [3, 45, 8, 55, 35, 34].
@@
Inspired by these advances, a new class of ML models, physics-informed machine learning, has gained traction within the earthquake engineering and structural dynamics community, where it addresses critical challenges such as data scarcity and limited model generalization. By embedding governing physical laws directly into ML architectures, physics-informed frameworks enhance predictive accuracy and ensure physical consistency. For instance, physics-informed LSTM frameworks (e.g., PhyLSTM2 and PhyLSTM3) [54] estimate key variables such as displacement and restoring force by incorporating these laws into the loss function during training. Similarly, PhyCNN [53] integrates domain knowledge with deep CNNs to predict the nonlinear seismic response of structures through a data-driven approach. These innovative methods have been applied across diverse scenarios, including subway station dynamics [15], vehicle–bridge interactions [28], and dam behavior prediction [29]. More recently, PI-LSTM has been introduced and benchmarked against PhyCNN and PhyLSTM, further demonstrating the potential of physics-guided ML architectures for structural response prediction [31]. Other works [38, 37, 44, 56] have highlighted the importance of incorporating physical constraints—such as ordinary and partial differential equations and boundary conditions—directly into deep learning models to improve stability and accuracy, even when data are limited.
Despite this progress, key limitations persist. LSTM and RNN models, while effective in modeling temporal dependencies, suffer from long training times and are prone to overfitting, particularly in low-data regimes [2, 46]. Deep CNN models, although powerful, require substantial computational resources when applied to long time series, and their performance degrades under large plastic deformations [22]. Moreover, capturing long-range interactions and nonlinearities in dynamic systems remains a challenge. To overcome these limitations and address the growing demand for reliable seismic response prediction, we propose a novel Physics-Informed U-Net–LSTM (PhyULSTM) framework. While U-Net architectures have been predominantly employed in image processing tasks such as medical imaging and semantic segmentation, we reformulate the U-Net into a causal one-dimensional architecture specifically designed for time-series modeling of structural response and ground motion signals.
The proposed 1D U-Net extracts hierarchical, multi-scale temporal representations while strictly preserving chronological causality—an essential requirement for structural health monitoring and early-warning applications. By leveraging the slow feature hypothesis [48], the encoder–decoder structure captures temporally coherent latent features that evolve smoothly across scales, enabling robust representation of nonlinear structural dynamics. These learned representations are subsequently processed through a deep LSTM network, which is excellent for modeling long-range temporal dependencies and complex hysteretic behavior inherent in nonlinear structural systems. Beyond purely data-driven training, physical knowledge in the form of governing equations of motion is embedded directly into the loss function via a physics-based residual term. This ensures mechanical consistency and enables accurate prediction even in data-scarce or partially observed scenarios. The framework is validated through both numerical and experimental benchmarks, where it consistently and significantly outperforms popular physics-informed models such as PhyLSTM and PhyCNN. Its superiority is firmly established through rigorous quantitative assessment, including high correlation coefficients between predicted and actual responses and superior global error metrics. Furthermore, detailed analysis of hysteresis plots, Fourier amplitude spectra, and peak response errors reveals that the model captures the system’s nonlinear behavior with exceptional fidelity, accurately reproduces critical frequency components, and precisely predicts peak responses—capabilities that are essential for subsequent fragility and reliability analysis of structures.
@@ 2 Physics-Informed U-net-LSTM Network
In this section, we present the PhyULSTM framework for surrogate modeling of nonlinear dynamic structural systems subjected to seismic ground motion. PhyULSTM consists of three core components: a 1D U-Net, a deep LSTM network, and a tensor differentiator. The proposed 1D U-Net, a tailored variant of the standard U-Net architecture, is designed to work synergistically with deep LSTM networks for time-series analysis. It has potential to alleviate common limitations of traditional RNNs and LSTMs—such as difficulty in capturing long-range temporal dependencies and susceptibility to overfitting—especially when processing high-resolution seismic inputs like ground acceleration. This is achieved by leveraging the ‘Slow Feature’ hypothesis [48], which posits that the most informative features in temporal sequences evolve slowly over time, even when raw signals exhibit rapid fluctuations. By extracting such temporally stable features across multiple time scales and applying causal convolutions—which restrict each prediction to depend only on current and past inputs—the U-Net enhances the LSTM’s ability to learn long-term dependencies within an auto-regressive framework. Additionally, physical knowledge is explicitly incorporated into the model through a tensor differentiator (based on the finite difference method), as introduced by Zhang et al. [53]. The differentiator component computes derivatives of the state-space outputs to construct a physics-informed loss term derived from the system’s governing equations. The hybrid U-Net–LSTM model is trained on seismic input–output datasets—obtained either through numerical simulation or sensor measurements—and subsequently deployed as a surrogate model for predicting structural responses under future seismic excitations. A detailed illustration of the PhyULSTM framework is presented in Figure 1.
@@
where $M$ is the mass matrix; $x(t)$ and $\dot{x}(t)$ denote the relative displacement and velocity with respect to the ground; $\ddot{x}_{g}(t)$ is the ground acceleration vector; $\Gamma$ represents the force distribution vector; and $h(t)$ denotes the latent generalized restoring force vector, which is modeled as a function of the ground acceleration $\ddot{x}_{g}(t)$ and the acceleration of structures with respect to the ground $\ddot{x}(t)$. Normalizing Eq. (1) by $M$, the governing equation can be rewritten as
@@
where $g(t)=M^{-1}h(t)$ is the mass-normalized restoring force.
Figure 1: The proposed Physics-Informed U-net-LSTM Network (PhyULSTM) for time-series modeling. The 1D U-net first receives the ground acceleration data as input and extracts features at multiple time scales. The outputs from the U-net are then fed to the deep LSTM network that maps the temporal feature maps to the corresponding output space. The outputs are state space variables z(t){z}(t) = $\{x(t),\dot{x}(t),g(t)\}$. Available physics knowledge is incorporated directly into the loss function. In addition to it, a tensor differentiator is implemented to calculate the derivative of state space outputs ${z_{t}}(t)$ = $\{x_{t}(t),\dot{x}_{t}(t),g_{t}(t)\}$ to construct the physics loss from the governing equation. By optimising the network hyperparameters, PhyULSTM can interpret the measurement data (e.g. $\{x_{m},\dot{x}_{m},g_{m}\}$) while satisfying the physical equation of motion in equation 2, e.g. $f\rightarrow 0$.
@@
Figure 2: Proposed 1D version of U-net: The U-Net architecture is mainly divided into four major components: the encoder blocks, bottleneck, decoder blocks, and the output convolution block. The encoder blocks perform convolution and pooling operations, generating skip connections and providing input to subsequent encoder blocks after pooling. The decoder blocks involve upsampling, followed by concatenation with the skip connections from the corresponding encoder block outputs, and subsequent convolution operations. The bottleneck layer, serving as a bridge, consists solely of a convolution block, linking the encoder and decoder paths. The output from the final decoder block is fed to the output convolution block, where it first passes through a Conv1D layer with a number of filters equal to the number of channels required in the final output, and sigmoid activation, followed by a second Conv1D layer and a linear activation function, which further processes the output to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension. Here, ‘ConvBlock’is a fundamental building block consisting of a convolutional layer, followed by Batch Normalisation, and a ReLU activation function.
@@
U-Net [40] is an established framework originally developed for biomedical image segmentation [42, 9]. It has a symmetrical U-shaped structure, consisting of an encoder and a decoder connected by skip connections. The encoder captures contextual information by progressively downsampling the input image, while the decoder reconstructs the image by upsampling and combining high-resolution features from the encoder, ensuring precise localisation. The ‘Slow Feature Analysis’ hypothesis [48] posits that meaningful features in dynamic signals often evolve slowly over time. Considering the encoder-decoder components and the ability to capture contextual information and combine resolution features, U-nets compute features at different time scales with two-dimensional convolutions and combining them to make predictions at the same resolution as the input. One such adaptation is the one-dimensional U-Net, suitable for time-series analysis [43]. In this study, we propose a causal 1D U-Net, as shown in Figure 2 and represented in the table 1, to process seismic time-series data that captures such features across multiple temporal scales while maintaining causal structure, ensuring outputs at a given time step depend only on current and past inputs. This enhances the model’s suitability for sequential tasks and integration with LSTM networks.
@@
The architecture comprises four parts: the encoder (contractive path), the bottleneck, the decoder part (expansive path) and finally, the output convolution block. The input is processed sequentially through two encoder blocks, having 50 and 100 filters respectively. Each encoder block generates two outputs: one is a skip connection (or shortcut connection) for the respective decoder block, and the other (output after 1D pooling operation) serves as the input for the subsequent encoder block. The final encoder block provides inputs for the bottleneck layer and also creates a skip connection. The encoder block comprises two sequential ‘ConvBlock’ units. Each ‘ConvBlock’ executes a series of operations: a 1D convolution ( kernel size = 2, padding = ‘causal’ ), followed by Batch Normalization, and finally a ReLU activation function. The resulting output from this process serves as both a skip connection (shortcut path) and as the input for the subsequent MaxPooling1D operation. MaxPooling1D is a down-sampling operation typically used in convolutional neural networks (CNNs). This method reduces the dimensionality of the input tensor by extracting the maximum value from within a sliding window, known as the pool size (set to 2), across the input. So, the encoder block consists of two main components: the convolution blocks and the max pooling block.
@@
For down-sampling, the ‘Conv1D’ operation with strides=2 (padding= ‘causal’) can also be implemented, but here in this case ‘MaxPooling1D’ is used to emphasize keeping feature extraction and down-sampling as distinct operations. And to ensure that the most significant features (e.g. sudden large significant peaks in ground acceleration data) are preserved during downsampling. There are also a few additional advantages. It is more straightforward than implementing ‘Conv1D’ with a stride of 2, and it reduces feature map size without adding extra learnable parameters. The bottleneck layer also comprises two sequential ‘ConvBlock’ units, using 200 filters without any pooling operation. This layer acts as a crucial bridge between the encoder and decoder, capturing complex, high-level features while maintaining the spatial dimensions of the feature maps. In the decoder stage, the first decoder block receives the output from the bottleneck layer, along with the skip connection from the second encoder block, and uses these inputs to upsample the feature map. Within the decoder block, a 1D upsampling operation is performed on the input tensor that repeats each temporal step size times along the time axis. In this case, as the parameter size=2 is specified, the layer performs upsampling by repeating each temporal step of the input sequence twice along the time axis. This step helps to gradually reconstruct the original input size. After the 1D upsampling operation, the feature map is concatenated with the corresponding skip connection from the encoder blocks. Then the output passes through two ConvBlocks. This sequence is repeated in each of the two decoder blocks, with 100 and 50 filters, respectively, to progressively refine the feature maps and reconstruct the final output with high accuracy and detail.
@@
The output from the final decoder block is then first passed through a Conv1D layer with a number of filters equal to the number of channels required in the final output, a kernel size of 1, and sigmoid activation, which reduces the channels to match the dimension of the final output required and maps the output values between 0 and 1. The output has the same length as the input due to ‘same’ padding. This is followed by a second Conv1D layer and a linear activation function, which further processes the output by expanding the channels to format the final output as a three-dimensional array, where the entries are sampled in the first dimension, time history steps in the second dimension, and output features in the last dimension, this is very crucial otherwise the deep LSTM network cannot be implemented subsequently. The network architecture is shown in Figure 2 and represented in the table 1. The U-Net output feeds directly into the LSTM module.
Table 1: Proposed 1D U-Net Architecture
Layer Operation Filters Kernel Stride Activation
@@
Output Convolutional $N_{\text{out}}$ 1 1 Sigmoid
Final Output Convolutional $N_{\text{out}}$ 1 1 Linear
Note: Each Conv Block consists of consecutive operations: Convolution $\rightarrow$ Batch Normalization $\rightarrow$ ReLU activation.